# CLIP-QDA: An Explainable Concept Bottleneck Model

### v1

@@ Abstract
In this paper, we introduce an explainable algorithm designed from a multi-modal foundation model, that performs fast and explainable image classification. Drawing inspiration from CLIP-based Concept Bottleneck Models (CBMs), our method creates a latent space where each neuron is linked to a specific word. Observing that this latent space can be modeled with simple distributions, we use a Mixture of Gaussians (MoG) formalism to enhance the interpretability of this latent space. Then, we introduce CLIP-QDA, a classifier that only uses statistical values to infer labels from the concepts. In addition, this formalism allows for both local and global explanations. These explanations come from the inner design of our architecture, our work is part of a new family of greybox models, combining performances of opaque foundation models and the interpretability of transparent models. Our empirical findings show that in instances where the MoG assumption holds, CLIP-QDA achieves similar accuracy with state-of-the-art methods CBMs. Our explanations compete with existing XAI methods while being faster to compute.
@@ 1 Introduction
  • We propose two efficient and mathematically grounded XAI metrics for model explanation. These metrics encompass both global and local assessments of why the model behaves. The global metric directly emanates from our Gaussian modeling approach, providing a comprehensive evaluation of CLIP-QDA’s performance. Additionally, our local metric draws inspiration from counterfactual analysis, furnishing insights into individual data points.
Figure 1: Overview of our modeling method. By considering the whole dataset CLIP scores ${\bm{z}}$ as class conditioned distributions ${\bm{Z}}=\begin{bmatrix}Z^{{\color[rgb]{0,0,0}1}}&\ldots&Z^{{\color[rgb]{0,0,0}N}}\end{bmatrix}$, we model the CLIP latent space as a mixture of Gaussians, allowing for mathematically grounded explanations.
@@ 2.1 Contrastive Image Language Pre-training (CLIP)
The widespread adoption of CLIP stems from the remarkable robustness exhibited by its pre-trained model. Through training on an extensive multimodal dataset, such as Schuhmann et al. (2022), the model achieves impressive performance. Thus, on few- and zero-shot learning, for which it was designed, it obtains impressive results across a wide range of datasets. Notably, CLIP provides a straightforward and efficient means of obtaining semantically rich representations of images in low-dimensional spaces. This capability enables researchers and practitioners to divert the original use of CLIP to various other applications (Luo et al., 2022; Menon & Vondrick, 2022).
While the concept of CBM has its origins prior to the rise of DNNs (Kumar et al., 2009; Lampert et al., 2009), and CLIP (Koh et al., 2020; Losch et al., 2019), the emergence of this multimodal foundational model has opened up novel opportunities. Recent research (Yang et al., 2023; Oikarinen et al., 2023) has leveraged large language models to directly construct concepts from CLIP text embeddings, opening the door to a family of CLIP-based CBMs. Additionally, efforts have been made to create sparse CLIP-based CBMs (Panousis et al., 2023; Feng et al., 2023). Yan et al. (2023a) explore methods to achieve superior representations with minimal labels. Yuksekgonul et al. (2022) capitalize on the CLIP embedding spaces, considering concepts as activation vectors. Finally, Kim et al. (2023) build upon the idea of activation vectors to discover counterfactuals.
@@ 2.2 Explainable AI
The literature shows a clear distinction between non-transparent (or blackbox) and transparent (or whitebox) models. Transparent models are characterized by their inherent explainability. These models can be readily explained due to their simplicity and easily interpretable features. Examples of such models include linear regression (Galton, 1886), logistic regression (McCullagh, 2019), and decision trees (Quinlan, 1986). In contrast, non-transparent models are inherently non-explainable. This category encompasses models that could have been explainable if they possessed simpler and more interpretable features (Galton, 1886; Quinlan, 1986), as well as models that inherently lack explainability, including deep neural networks. The distinction between these two types of models highlights the trade-off between model complexity and interpretability (Arrieta et al., 2020), with transparent models offering inherent explainability while non-transparent models allow for better performance, but require the use of additional techniques for explanations, named post-hoc methods. Commonly used post-hoc methods include visualization techniques, such as saliency maps (Selvaraju et al., 2017), which highlight the influential features in an image that contribute to the model’s decision-making. Sensitivity analysis (Cortez & Embrechts, 2011) represents another avenue, involving the analysis of model predictions by varying input data. Local explanation techniques are also used to explain the model from a local simplification of the model around a point of interest (Ribeiro et al., 2016; Plumb et al., 2018). Finally, feature relevance techniques aim at estimating the impact of each feature on the decision (Lundberg & Lee, 2017).
@@
3 Method
@@ 3.1 General framework
The selection of concepts is guided by expert input and acts as a hyperparameter within our framework. For comprehensive examples of concept sets, please refer to Section A.1. It is noteworthy that there is no requirement for individual image annotation with these concepts. This is due to CLIP’s inherent design, which allows it to score concepts in a zero-shot manner.
@@ 3.2.1 Notations and formalism
Let $\{{\bm{k}}^{j}\}_{j=1}^{N}$ denote a set of $N\in\mathbb{N}$ concepts, where each ${\bm{k}}^{j}$ is a character string representing the concept in natural language. We consider “CLIP’s DNN” to refer to the vector of its pre-trained weights, denoted as ${\bm{\omega}}^{g}$, and a function $g$ that represents the architecture of the deep neural network (DNN). Given an image ${\bm{x}}$ and a concept ${\bm{k}}^{j}$, the output of CLIP’s DNN is represented as $z^{j}=g_{{\bm{\omega}}^{g}}({\bm{x}},{\bm{k}}^{j})$. The projection in the multi-modal latent space of an image ${\bm{x}}$ is the vector ${\bm{z}}=\begin{bmatrix}g_{{\bm{\omega}}^{g}}({\bm{x}},{\bm{k}}^{{\color[rgb]{0,0,0}1}})&\ldots&g_{{\bm{\omega}}^{g}}({\bm{x}},{\bm{k}}^{{\color[rgb]{0,0,0}N}})\end{bmatrix}$. We define $Z^{j}$ as the random variable associated with the observation $z^{j}$. It should be noted that ${\bm{Z}}=\begin{bmatrix}Z^{{\color[rgb]{0,0,0}1}}&\ldots&Z^{{\color[rgb]{0,0,0}N}}\end{bmatrix}$ is the random variable representing the concatenation of the CLIP scores associated with the $N$ concepts. Furthermore, we denote the conditional distributions of ${\bm{z}}$ having class $c$ as ${\bm{Z}}_{Y=c}=\begin{bmatrix}Z^{{\color[rgb]{0,0,0}1}}_{Y=c}&\ldots&Z^{{\color[rgb]{0,0,0}N}}_{Y=c}\end{bmatrix}$.
Finally, we define the classifier as a function $h_{{\bm{\omega}}^{h}}({\bm{z}})$ with parameters ${\bm{\omega}}^{h}$ that, given a vector ${\bm{z}}$, attributes the predicted class.
@@ 3.2.2 Gaussian modeling of CLIP’s latent space
$p({\bm{Z}}={\bm{z}})=$ $\sum_{c=1}^{C}p_{c}\mathcal{N}({\bm{z}}\mid{\bm{\mu}}_{c},{\bm{\Sigma}}_{c})\ .$ (2)
3.3 CLIP Quadratic Discriminant Analysis
@@
$p(Y=c\mid{\bm{Z}}={\bm{z}})=\frac{p_{c}\mathcal{N}({\bm{z}}\mid{\bm{\mu}}_{c},{\bm{\Sigma}}_{c})}{\sum_{c=1}^{N}p_{c}\mathcal{N}({\bm{z}}\mid{\bm{\mu}}_{c},{\bm{\Sigma}}_{c})}\,.$ (3)
@@
3.4 XAI Metrics
Since our description of the CLIP latent space is grounded in a probabilistic framework, we have access to a range of statistical tools for interpreting the way our classifier works. In this section, we introduce two distinct types of explanations: a global explanation that provides insights into the way the classifier acts on the entire dataset (see Section 3.4.1), and a local explanation designed to elucidate the model’s behavior on a specific sample (see Section 3.4.2).
3.4.1 Global metric
@@
To accomplish this, we propose to use the Wasserstein-2 distance as a metric for quantifying the separation between the two conditional distributions. It is worth noting that calculating the Wasserstein-2 distance can be a complex task in general. However, for Gaussian distributions, there exists a closed-form solution for computing the Wasserstein-2 distance. In addition, we sign the distance to keep the information of the position of $c_{1}$ relative to $c_{2}$:
@@
where $\Lambda^{j}_{c_{1},c_{2}}=[{\bm{\Sigma}}_{c_{1}}]_{(j,j)}+[{\bm{\Sigma}}_{c_{2}}]_{(j,j)}-2{\color[rgb]{0,0,0}\sqrt{[{\bm{\Sigma}}_{c_{1}}]_{(j,j)}[{\bm{\Sigma}}_{c_{2}}]_{(j,j)}}}$.
Note that the resulting value is no longer a distance since we lost the commutativity property. Examples of explanations based on this metric are given in Sections 4.4 and A.4.
3.4.2 Local metric
@@
Consider a pre-trained classifier denoted as $h_{{\bm{\omega}}^{h}}(\cdot)$. In this context, ${\bm{\omega}}^{h}$ represents the set of weights associated with the CLIP-QDA, specifically ${\bm{\omega}}^{h}=({\bm{\Sigma}}_{c},{\bm{\mu}}_{c},p_{c})$. Given a score vector ${\bm{z}}$, we define counterfactuals as hypothetical values ${\bm{z}}+{\bm{\epsilon}}^{j}_{{\color[rgb]{0,0,0}s}}$, ${\bm{\epsilon}}^{j}_{{\color[rgb]{0,0,0}s}}$ being called the perturbation. This perturbation aims to be of minimal magnitude and is obtained by solving the following optimization problem:
$\min\|{\bm{\epsilon}}^{j}_{{\color[rgb]{0,0,0}s}}\|^{2}~{}~{}\textnormal{s.t.}~{}~{}h_{{\bm{\omega}}^{h}}({\bm{z}}+\epsilon^{j}_{s})\neq h_{{\bm{\omega}}^{h}}({\bm{z}})\,.$ (5)
The idea behind this equation is to find the minimal perturbation ${\bm{\epsilon}}^{j}_{{\color[rgb]{0,0,0}s}}$ of the input ${\bm{z}}$ that makes the classifier produce a different label than $h_{{\bm{\omega}}^{h}}({\bm{z}})$. However, in our case, two important restrictions are applied to ${\bm{\epsilon}}^{j}_{{\color[rgb]{0,0,0}s}}$:
@@
    Sparsity: for interpretability, we only change one attribute at a time, indicated by the index $j$. Then ${\bm{\epsilon}}^{j}_{{\color[rgb]{0,0,0}s}}=[0,..,0,\epsilon^{j}_{{\color[rgb]{0,0,0}s}},0,...,0]$.
@@
    Sign: we take into account the sign $s\in\{-,+\}$ of the perturbation. Then, we separate the positive counterfactuals ${\bm{\epsilon}}^{j}_{{\color[rgb]{0,0,0}+}}=[0,..,0,\epsilon^{j}_{+},0,...,0],~{}\epsilon^{j}_{+}\in\mathbb{R^{+}}$ and the negative counterfactuals. ${\bm{\epsilon}}^{j}_{-}=[0,..,0,\epsilon^{j}_{-},0,...,0],~{}\epsilon^{j}_{-}\in\mathbb{R^{-}}$.
@@
The proof and expression are given in Appendix A.2. We illustrate the behavior of our classifier and our local metric with a toy example which consists of two Gaussians ($C=2$) among two concepts $Z_{1}$ and $Z_{2}$ ($N=2$). We find the counterfactuals for both signs following the first concept (${\bm{\epsilon}}^{{\color[rgb]{0,0,0}1}}_{-,*}$ and ${\bm{\epsilon}}^{{\color[rgb]{0,0,0}1}}_{+,*}$). Results are presented in Figure 4.
@@
Then, the value of each counterfactual can be expressed as “the addition (or subtraction) of standard deviations in accordance to ${\bm{Z}}_{Y=h_{{\bm{\omega}}^{h}}({\bm{z}})}$ that changes the label”. Examples of such explanations are given in Sections 4.4 and A.4.

### v2

@@ Abstract
In this paper, we introduce an explainable algorithm designed from a multi-modal foundation model, that performs fast and explainable image classification. Drawing inspiration from CLIP-based Concept Bottleneck Models (CBMs), our method creates a latent space where each neuron is linked to a specific word. Observing that this latent space can be modeled with simple distributions, we use a Mixture of Gaussians (MoG) formalism to enhance the interpretability of this latent space. Then, we introduce CLIP-QDA, a classifier that only uses statistical values to infer labels from the concepts. In addition, this formalism allows for both sample-wise and dataset-wise explanations. These explanations come from the inner design of our architecture, our work is part of a new family of greybox models, combining performances of opaque foundation models and the interpretability of transparent models. Our empirical findings show that in instances where the MoG assumption holds, CLIP-QDA achieves similar accuracy with state-of-the-art methods CBMs. Our explanations compete with existing XAI methods while being faster to compute.
@@ 1 Introduction
  • We propose two efficient and mathematically grounded XAI methods for model explanation, named CLIP-QDAlocal and CLIP-QDAglobal. These methods encompass both global and local assessments of why the model behaves. The global metric directly emanates from our Gaussian modeling approach, providing a comprehensive evaluation of CLIP-QDA’s performance. Additionally, our local metric draws inspiration from counterfactual analysis, furnishing insights into individual data points.
  • We extend two established post-hoc XAI methods, LIME and SHAP, to formulate a novel XAI approach specifically tailored for CBMs. Departing from the conventional application of these methods, which typically produce explanations on the image level, CLIP-LIME and CLIP-SHAP generate explanations on the concept level allowed by CBMs.
  • We propose a new evaluation protocol,specifically designed for the unique characteristics of CBMs to assess the effectiveness of explanations. This protocol includes a deletion metric, which examines faithfulness to the model, and a detection metric, which evaluates faithfulness to the data.
Figure 1: Overview of our modeling method. By considering the whole dataset CLIP scores ${\bm{z}}$ as class conditioned distributions ${\bm{Z}}=\begin{bmatrix}Z^{{1}}&\ldots&Z^{{N}}\end{bmatrix}$, we model the CLIP latent space as a mixture of Gaussians, allowing for mathematically grounded explanations.
@@ 2.1 Contrastive Image Language Pre-training (CLIP)
The widespread adoption of CLIP stems from the remarkable robustness exhibited by its pre-trained model. Through training on an extensive multimodal dataset, such as Schuhmann et al. (2022), the model achieves impressive performance. Thus, on few- and zero-shot learning, for which it was designed, it obtains impressive results across a wide range of datasets. Notably, CLIP provides a straightforward and efficient means of obtaining semantically rich representations of images in low-dimensional spaces. This capability enables researchers and practitioners to divert the original use of CLIP to various other applications (Luo et al., 2022; Menon & Vondrick, 2022; Gabeff et al., 2023).
2.2 CLIP-based Concept Bottleneck Models (CLIP-based CBMs)
The term Concept Bottleneck Model (CBM), as outlined in Koh et al. (2020), pertains to the implementation of a bottleneck reliant on human-specified concepts to execute a task, predominantly image classification. Consequently, the resultant algorithm inherently facilitates enhanced interpretability. While the term itself is relatively recent, it characterizes a lineage of methods that were employed in earlier research (Kumar et al., 2009; Lampert et al., 2009; Koh et al., 2020; Losch et al., 2019). However, despite the advantages offered by CBMs in terms of better understanding, early implementations encountered challenges stemming from the requirement for dedicated datasets. These datasets needed to encompass not only inputs and labels but also incorporate human-specified concepts for each sample.
In this context, the emergence of multimodal foundational models has opened up novel opportunities. Recent research (Yang et al., 2023; Oikarinen et al., 2023) has leveraged large language models to directly construct concepts from CLIP text embeddings, opening the door to a family of CLIP-based CBMs. Additionally, efforts have been made to create sparse CLIP-based CBMs (Panousis et al., 2023; Feng et al., 2023). Yan et al. (2023a) explore methods to achieve superior representations with minimal labels. Yuksekgonul et al. (2022) capitalize on the CLIP embedding spaces, considering concepts as activation vectors. Finally, Kim et al. (2023) build upon the idea of activation vectors to discover counterfactuals.
@@ 2.3 Explainable AI
The literature shows a clear distinction between non-transparent (or blackbox) and transparent (or whitebox) models. Transparent models are characterized by their inherent explainability. These models can be readily explained due to their simplicity and easily interpretable features. Examples of such models include linear regression (Galton, 1886), logistic regression (McCullagh, 2019), and decision trees (Quinlan, 1986). In contrast, non-transparent models are inherently non-explainable. This category encompasses models that could have been explainable if they possessed simpler and more interpretable features (Galton, 1886; Quinlan, 1986), as well as models that inherently lack explainability, including deep neural networks. The distinction between these two types of models highlights the trade-off between model complexity and interpretability (Arrieta et al., 2020), with transparent models offering inherent explainability while non-transparent models allow for better performance, but require the use of additional techniques for explanations, named post-hoc methods. Commonly used post-hoc methods include visualization techniques, such as saliency maps, which highlight the influential features in an image that contribute to the model’s decision-making. Within this category of methods, notable examples include approaches such as Grad-CAM (Selvaraju et al., 2017), which generates activation maps by computing the gradients of the output labels. Sensitivity analysis (Cortez & Embrechts, 2011) represents another avenue, involving the analysis of model predictions by varying input data. Sample-wise explanation techniques are also used to explain the model from a local simplification of the model around a point of interest (Ribeiro et al., 2016; Plumb et al., 2018). Finally, feature relevance techniques aim at estimating the impact of each feature on the decision (Lundberg & Lee, 2017).
@@
Feature Attribution Methods, a category of techniques employed to address the complexity of DNNs’ output for explainability, strive to identify crucial features in the input. These methods leverage a mapping function to reduce input complexity. Notable examples in this family include DeepSHAP and KernelSHAP, as proposed by Lundberg & Lee (2017). This approach resonates with the concept of greybox models, wherein the input is initially simplified to be explainable by a transparent classifier. However, greyboxes differ from Feature Attribution Methods in that the mapping function is independent from the input under consideration.
3 A Greybox Concept Bottleneck Model: CLIP-QDA
@@ 3.1 General Framework
The selection of concepts is guided by expert input and acts as a hyperparameter within our framework. For comprehensive examples of concept sets, please refer to Section A.3. Notably, there is no requirement for individual image annotation with these concepts. This is due to CLIP’s inherent design, which allows it to score concepts in a zero-shot manner.
@@ 3.2.1 Notations and formalism
Let $\{{\bm{k}}^{j}\}_{j=1}^{N}$ denote a set of $N\in\mathbb{N}$ concepts, where each ${\bm{k}}^{j}$ is a character string representing the concept in natural language. We consider “CLIP’s DNN” to refer to the vector of its pre-trained weights, denoted as ${\bm{\omega}}^{g}$, and a function $g$ that represents the architecture of the deep neural network (DNN). Given an image ${\bm{x}}$ and a concept ${\bm{k}}^{j}$, the output of CLIP’s DNN is represented as $z^{j}=g_{{\bm{\omega}}^{g}}({\bm{x}},{\bm{k}}^{j})$. The projection in the multi-modal latent space of an image ${\bm{x}}$ is the vector ${\bm{z}}=\begin{bmatrix}g_{{\bm{\omega}}^{g}}({\bm{x}},{\bm{k}}^{{1}})&\ldots&g_{{\bm{\omega}}^{g}}({\bm{x}},{\bm{k}}^{{N}})\end{bmatrix}$. We define $Z^{j}$ as the random variable associated with the observation $z^{j}$. It should be noted that ${\bm{Z}}=\begin{bmatrix}Z^{{1}}&\ldots&Z^{{N}}\end{bmatrix}$ is the random variable representing the concatenation of the CLIP scores associated with the $N$ concepts. Furthermore, we denote the conditional distributions of ${\bm{z}}$ having class $c$ as ${\bm{Z}}_{Y=c}=\begin{bmatrix}Z^{{1}}_{Y=c}&\ldots&Z^{{N}}_{Y=c}\end{bmatrix}$.
Finally, we define the classifier as a function $h_{{\bm{\omega}}^{h}}({\bm{z}})$ with parameters ${\bm{\omega}}^{h}$ that, given a vector ${\bm{z}}$, concepts the predicted class.
@@ 3.2.2 Gaussian modeling of CLIP’s latent space
$p({\bm{Z}}={\bm{z}})=$ $\sum_{{c_{i}}=1}^{C}p_{{c_{i}}}\mathcal{N}({\bm{z}}\mid{\bm{\mu}}_{{c_{i}}},{\bm{\Sigma}}_{{c_{i}}})\ .$ (2)
3.3 CLIP Quadratic Discriminant Analysis (CLIP-QDA)
@@
$p(Y=c\mid{\bm{Z}}={\bm{z}})=\frac{p_{c}\mathcal{N}({\bm{z}}\mid{\bm{\mu}}_{c},{\bm{\Sigma}}_{c})}{\sum_{{c_{i}}=1}^{N}p_{{c_{i}}}\mathcal{N}({\bm{z}}\mid{\bm{\mu}}_{{c_{i}}},{\bm{\Sigma}}_{{c_{i}}})}\,.$ (3)
@@
3.4 Explainable AI for Concept Bottleneck Models
Our CLIP-QDA model is founded upon a transparent probabilistic framework, hence, we have at our disposal a variety of statistical tools to explain the functioning of our classifier as illustrated in Figure 4. In this section, we present two distinct types of explanations: CLIP-QDAglobal, offering a global perspective that sheds light on the classifier’s behavior across the entire dataset (refer to Section 3.4.1), and CLIP-QDAlocal, providing a local explanation tailored to elucidate the model’s actions on individual samples (refer to Section 3.4.2).
Furthermore, we propose the adaptation of two well-established XAI post-hoc methods, LIME (Ribeiro et al., 2016) and SHAP (Lundberg & Lee, 2017), to the unique characteristics of CBMs. These methods are denoted as CLIP-SHAP and CLIP-LIME, respectively (see Section 3.4.3). An overview of the methods incorporated in our investigation is depicted in Figure 4.
Figure 4: The different source of explanation of CLIP-QDA. Global (CLIP-QDAglobal) vs. Local (CLIP-QDAlocal) explanations, offering insights into classifier behavior across the dataset and on individual samples, respectively. Post-hoc explanations with CLIP-SHAP and CLIP-LIME offering using traditional techniques.
3.4.1 Dataset-wise explanation with CLIP-QDAglobal
@@
To accomplish this, we propose to use the Wasserstein-2 distance (Ramdas et al., 2017) as a metric for quantifying the separation between the two conditional distributions. It is worth noting that calculating the Wasserstein-2 distance can be a complex task in general. However, for Gaussian distributions, there exists a closed-form solution for computing the Wasserstein-2 distance. In addition, we sign the distance to keep the information of the position of $c_{1}$ relative to $c_{2}$:
@@
where $\Lambda^{j}_{c_{1},c_{2}}=[{\bm{\Sigma}}_{c_{1}}]_{(j,j)}+[{\bm{\Sigma}}_{c_{2}}]_{(j,j)}-2{\sqrt{[{\bm{\Sigma}}_{c_{1}}]_{(j,j)}[{\bm{\Sigma}}_{c_{2}}]_{(j,j)}}}$.
Note that the resulting value is no longer a distance since we lost the commutativity property. Examples of explanations based on this metric are given in Sections 4.5 and A.7. Also, an alternative way to produce dataset-wise explanations, oriented on example based explanations, are also available in Section A.4.
3.4.2 Sample-wise explanation with CLIP-QDAlocal
@@
Consider a pre-trained classifier denoted as $h_{{\bm{\omega}}^{h}}(\cdot)$. In this context, ${\bm{\omega}}^{h}$ represents the set of weights associated with the CLIP-QDA, specifically ${\bm{\omega}}^{h}=({\bm{\Sigma}}_{c},{\bm{\mu}}_{c},p_{c})$. Given a score vector ${\bm{z}}$, we define counterfactuals as hypothetical values ${\bm{z}}+{\bm{\epsilon}}^{j}_{{s}}$, ${\bm{\epsilon}}^{j}_{{s}}$ being called the perturbation. This perturbation aims to be of minimal magnitude and is obtained by solving the following optimization problem:
$\min\|{\bm{\epsilon}}^{j}_{{s}}\|^{2}~{}~{}\textnormal{s.t.}~{}~{}h_{{\bm{\omega}}^{h}}({\bm{z}}+\epsilon^{j}_{s})\neq h_{{\bm{\omega}}^{h}}({\bm{z}})\,.$ (5)
The idea behind this equation is to find the minimal perturbation ${\bm{\epsilon}}^{j}_{{s}}$ of the input ${\bm{z}}$ that makes the classifier produce a different label than $h_{{\bm{\omega}}^{h}}({\bm{z}})$. However, in our case, two important restrictions are applied to ${\bm{\epsilon}}^{j}_{{s}}$:
@@
    Sparsity: for interpretability, we only change one attribute at a time, indicated by the index $j$. Then ${\bm{\epsilon}}^{j}_{{s}}=[0,..,0,\epsilon^{j}_{{s}},0,...,0]$.
@@
    Sign: we take into account the sign $s\in\{-,+\}$ of the perturbation. Then, we separate the positive counterfactuals ${\bm{\epsilon}}^{j}_{{+}}=[0,..,0,\epsilon^{j}_{+},0,...,0],~{}\epsilon^{j}_{+}\in\mathbb{R^{+}}$ and the negative counterfactuals. ${\bm{\epsilon}}^{j}_{-}=[0,..,0,\epsilon^{j}_{-},0,...,0],~{}\epsilon^{j}_{-}\in\mathbb{R^{-}}$.
@@
The proof and expression are given in Appendix A.5. We illustrate the behavior of our classifier and our local metric with a toy example which consists of two Gaussians ($C=2$) among two concepts $Z_{1}$ and $Z_{2}$ ($N=2$). We find the counterfactuals for both signs following the first concept (${\bm{\epsilon}}^{{1}}_{-,*}$ and ${\bm{\epsilon}}^{{1}}_{+,*}$). Results are presented in Figure 5.
@@
Then, the value of each counterfactual can be expressed as “the addition (or subtraction) of standard deviations in accordance to ${\bm{Z}}_{Y=h_{{\bm{\omega}}^{h}}({\bm{z}})}$ that changes the label”. Examples of such explanations are given in Sections 4.5 and A.7.
3.4.3 CLIP-LIME and CLIP-SHAP
CLIP-LIME.
To adapt LIME to the operation of CBMs, we begin with the image input $x$. From this input, we calculate the projection in the latent space ${\bm{z}}=\begin{bmatrix}g_{{\bm{\omega}}^{g}}({\bm{x}},{\bm{k}}^{{1}})&\ldots&g_{{\bm{\omega}}^{g}}({\bm{x}},{\bm{k}}^{{N}})\end{bmatrix}$. Subsequently, following the LIME method, we train a surrogate model to approximate $h_{{\bm{\omega}}^{h}}$ in the vicinity of ${\bm{z}}$ by training it on a dataset comprised of perturbed inputs around ${\bm{z}}$. Finally, the explanation is derived from the importance weights of the resulting surrogate model.
CLIP-SHAP.
To adapt SHAP to CBMs, we also consider the projection ${\bm{z}}=\begin{bmatrix}g_{{\bm{\omega}}^{g}}({\bm{x}},{\bm{k}}^{{1}})&\ldots&g_{{\bm{\omega}}^{g}}({\bm{x}},{\bm{k}}^{{N}})\end{bmatrix}$. To compute statistical values relevant to CLIP-SHAP, we incorporate the projections of all the images in the training set. Given that the number of concepts can be high, we employ the Kernel version of SHAP (Kernel SHAP) to reduce the computational cost. The resulting explanation comprises the Shapley values associated with each of the concepts.